OpenStudy (anonymous):

how do you figure out the domain of an equation?? ex. (x^5 - 2)^7 .. what is the domain? .. but i want to know how to solve on my own

OpenStudy (anonymous):

domain is the values x can be. so, for instance, if you have an equation like y = x, x could be any number, there is nothing restricting it's value

OpenStudy (anonymous):

if you have an equation like y = 1/x, x cannot be zero because you cannot divide by zero (it is undefined, doesn't make sense to split things up into 'zero' parts)

OpenStudy (anonymous):

So, in essence, the domain is all of the possible values you can put into an equation. On the flip side, the range is all the values that can come out. (irrelevant but good to know)

OpenStudy (anonymous):

ok .. so any equation with a fraction that has a zero under is undefined? right

OpenStudy (anonymous):

Well, it is undefined when x = 0. But it is defined for all other numbers other than 0.

OpenStudy (anonymous):

aka the domain is every real number except 0

OpenStudy (anonymous):

there are also several different ways to write the domain one way to state the domain of f(x) = 1/x is -infinity < 0 < infinity meaning x can be equal to any number besides 0

OpenStudy (anonymous):

ok so the equation i showed.. since its all real numbers .. does that mean the answer would be \[(\infty, -\infty)\]

OpenStudy (anonymous):

well, thats not really a good way to write it, but i think you get the idea

OpenStudy (anonymous):

how would i write it?

OpenStudy (anonymous):

the usual way to write it would be (-inf,inf)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

yeh.. i think im getting the hang of this... thank you so much!!!!

OpenStudy (anonymous):

but its good to learn the domain of all basic functions all polynomials have (-inf,inf) as a domain. the equation you listed is a polynomial, thus, it has that domain.

OpenStudy (anonymous):

oh ok... im sorry but please explain further.. or give an example

OpenStudy (anonymous):

look there at the first table. it will give you the domain of all basic functions

OpenStudy (anonymous):

the first 5 cases listed are all polynomial equations

OpenStudy (anonymous):

the first being a linear equation, 2nd quadratic, 3rd cubic, and then the general case x^n

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

basically x raised to any number will always have (-inf,inf) has its domain

OpenStudy (anonymous):

x^2 (x-2)^2 4x + 3x^2 - (x+4)^2 + x^5 all of these are polynomial equations, all have (-inf,inf) for a domain

OpenStudy (anonymous):

if you understand the domain/range of the functions you have to use (there are only 8-10 ones you really need to know) it makes algebra a whole lot easier

OpenStudy (anonymous):

i see that it says x^2 has -inf,inf and range is 0,inf and then it says f(x)=square root x domain and range is 0,inf why? arent they the same?

OpenStudy (anonymous):

that is a special case where the power you are raising to is less than 0 you must understand that exponentiation is repeated multiplication like, 2^4 just means 2*2*2*2 (2 multiplied by itself 4 times)

OpenStudy (anonymous):

so what does it mean when you have something like 2^(1/2)? you are multiplying 2 by itself... 1/2 times?

OpenStudy (anonymous):

or in other words, what number multiplied by itself will give you 2?

OpenStudy (anonymous):

the point of what i'm saying, is the domain of something like x^1/2 is [0,inf)

OpenStudy (anonymous):

the reason is because what you are doing is finding a number such that when you multiply it by itself it equals x

OpenStudy (anonymous):

now the problem is, when you multiply a number by itself, it can never be negative

OpenStudy (anonymous):

that is, a*a will always be a^2 and -a*-a will always be a^2

OpenStudy (anonymous):

thus, a^2 can never be negative

OpenStudy (anonymous):

thus, the domain of something like x^1/2 can never be negative

OpenStudy (anonymous):

i hope that makes sense, it's kind've complicated with words lol

OpenStudy (anonymous):

sorry.. im folling .. just got confused by the number that you get multiplied by itself to get 2

OpenStudy (anonymous):

like, what is the square root of 4?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

yes, so the answer to a square root problem looks like this sqrt(4) = 2*2

OpenStudy (anonymous):

in other words, the answer to a square root problem is a number multiplied by itself (in this case, 2)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yes, alright : ) now you see, the square root of 4 can never be negative

OpenStudy (anonymous):

because 2*2 = 4 and (-2)*(-2) = 4 also

OpenStudy (anonymous):

oooooooohhh!!!!

OpenStudy (anonymous):

ok i got it lol

OpenStudy (anonymous):

so, the square root function can never have a domain that is negative... simply beacuse it cant ever get there!

OpenStudy (anonymous):

alright cool haha hope that helps try to look at the other basic functions and figure them out if you want. it will help to understand problems a loooot

OpenStudy (anonymous):

i would also like to point out, that any time the power is negative like x^3 or x^5 or x^7 etc.

OpenStudy (anonymous):

then, the answer CAN be negative

OpenStudy (anonymous):

or woops what i mean is x^1/3 x ^1/5 x^1/7

OpenStudy (anonymous):

if the fraction is an odd number (like 3,5,7) then you can get negatives because (-a)*(-a)*(-a) is actually -(a^3)

OpenStudy (anonymous):

in other words, if you multiply a number by itself 3 times, it can be negative in like in the case of only 2 numbers, which cannot be negative.

OpenStudy (anonymous):

hope you understand that, gotta go eat some food enjoy

OpenStudy (anonymous):

ok.. thanks .. i think i got it down!! :D .. incase of anything ill ask more questions later.. thank you so much again!!! enjoy getting food!!! have a feast!!! lol